Ralphs age less 10 is equal to half of the sum of his age and 26
Answer by Shana-D77(132)



Usually we see "less than" as in "7 less than 10", which is 10-7 = 3.
In your case, you have just "less", which is way old school.
Example: "7 less 10" is 7 - 10 = -3


Let Ralph's age = R
"Ralphs age less 10"
R - 10
"is equal to"
=
"half of"
1/2 times
"the sum of his age and 26"
(R + 26)
Let's put it together:
R - 10 = 1/2(R + 26)
Let's clear the fraction by multiplying everything my 2 (to get rid of the 1/2)
2R - 20 = R + 26
R - 20 = 26 (subtracted R from both sides)
R = 46
So, Ralph is 46 years old